Given the equation of a sphere "x^{2}+y^{2}+z^{2}-4x+2y+2z=-5"

Which of the following is the vector equation of the sphere above?



  1. "|\mathbf{r}-(2,1,1)|=1"
  2. "|\mathbf{r}-(2,-1,-1)|=1"
  3. "|\mathbf{r}-(2,-1,1)|=1"
  4. "\u2223r\u2212(2,\u22121,1)\u2223=1."
  5. "\u2223r\u2212(2,1,\u22121)\u2223=1."
1
Expert's answer
2021-10-21T14:56:33-0400

Given the equation of the sphere

"x^2+y^2+z^2-yx+2y+2z=-5\\\implies xx^2-4x+y^2+2y+z^2+2z=-5\\\implies (x^2-4x+2^2-2^2)+(y^2+2y+1-1)+(z^2+2Z+1-1)=-5\\ [(x-2)^2-4]+[(y+1)^2-1]+[(z+1)^2-1]=-5\\\implies (x-2)^2+(y+1)^2+(z+1)^2-6=-5\\\implies (x-2)^2+(y+1)^2+(z+1)^2=1\\"


centre of sphere "=(2,-1,-1)"


radius of sphere "=1"


vector form is "|r-(2,-1,-1)|=1"


Hence option 2 is correct